# MAVIL: Design of a Multidimensional Assessment of Visual Data Literacy and its Application in a Representative Survey

### v2

@@ 3 MAVIL: A Multidimensional VDL Assessment
Dimension 1: Comprehending. First, self-perception of overall graph comprehension, including abstract thinking, is surveyed. The general impressions of the provided data visualization example are addressed (Q1a/Q1b). Further, following a definition of deductive and inductive abstract thinking as an indicator of a high ability level [7], respondents are asked if they could easily summarize the content of the visualization to someone (Q1c).
@@
Dimension 2: Decoding. Dimension 2 focuses on self-perceived decoding of graph specifics, such as graph elements and familiarity with chart types. The comprehensibility of color usage (Q2b), like the comprehensibility of depicted measurement units (Q2a), is queried. This reflects the understanding that visualization comprehension depends on both structural/content-related features and visual/aesthetic design elements [70]. Moreover, graph familiarity is said to be beneficial for reading data correctly [31], and therefore also investigated (Q2c).
@@
Dimension 3: Aestheticizing. Dimension 3 addresses self-assessed aesthetic perception in relation to Visual Data Literacy. Here, the BeauVis scale [41] was adapted (Q3a-Q3c). This scale consists of singular, English-language adjectives, all synonymous and intended to iterate on aesthetic perception. We translated BeauVis’ descriptive terms into German, the primary language of the surveyed country, which was done in a language-sensitive manner to convey their initial idea. The questions on aesthetic perception are supplemented with a question on perceived trustworthiness (Q3d), as attributing beauty to a visualization may favorably bias its interpretation [18].
Figure 1: Structure of the MAVIL assessment. Questions are grouped into six Visual Data Literacy dimensions (Dim1–Dim6), alongside demographic questions. Question formats include Likert scale ratings, text fields, true/false tasks, and multiple-choice items. For the full question catalog, see the supplementary material.
@@
Figure 2: Data visualizations used in the MAVIL survey with a German-speaking audience. Figure (a) shows an official English version of the tested line chart, while Figure (b) displays the orginial bar chart used in the survey, which is only available in German. See Section 4.1.1 for details.
Dimension 4: Critiquing. Dimension 4 centers on critical ability and understanding of visualization construction. Respondents are asked to articulate specific feedback on the data visualization. This includes indicating whether they have any erroneous findings to report (Q4a) or specific amendment proposals to share (Q4b/Q4c), as literacy covers the ability to question an object and to seek out possible changes [17]. This also plays into the idea of Visual Literacy as a grouping of acquired abilities and skills covering not only reading but also creating visualizations [10]. Respondents were offered subsequent optional text fields asking for visualization strengths and weaknesses (Q4d/Q4e).
Dimension 5: Reading. Dimension 5 complements MAVIL with task-based assessment questions by presenting data reading exercises (Q5a-Q5d), as this type of assessment question is a well established performance measure for literacy (see Section 2.2). Similar to related styles of assessments, such as in VLAT [78], these tasks directly refer to the data content of the surveyed visualization, as exemplified through true-or-false questions implemented for our surveyed visualizations in Figure 2.
Dimension 6: Contextualizing. Finally, two contextual ability facets closely related to Visual Data Literacy are surveyed: self-perception of numeracy skills and topic knowledge. A higher level of numeracy would indicate a higher likeliness of reading information from a data visualization correctly [69, 81], which is why we asked respondents rate their own numeracy level (Q6b). Beyond addressing this particular skill, personal interest in the displayed content is questioned (Q6a), as it may influence the understanding of visually represented data [31].
Even though most questions are generally phrased, some reference the data visualization example directly to exemplify parts of the questions to respondents. This is the case for data reading tasks (Q5a-Q5d) bound to the visualization content. Specifications were also made to Q2a by listing measurement units and Q2c by naming the chart type. In Q6a, respondents were asked to rate personal interest in climate-related topics.
@@ 4 Methodology
The design, development, and testing of MAVIL was split into four phases. At first, questions were developed in a theory-driven approach (for detail see Section 3) and placed within a digital survey format. Then, in two phases of qualitative testing, the assessment was presented to eight participants. Semi-structured interviews allowed us to gain insights into the validity of MAVIL questions, which were followed by iterative development. Finally, the assessment was completed by 438 respondents representing a German-speaking country’s age groups, from 18 to 74 years, and male-female gender distribution. MAVIL was developed and tested in the country’s primary language. The English version presented here is a translation and has not been separately tested. The original questions are made available at an anonymized link (to be provided in the camera-ready version).
4.1 MAVIL in Practice: Designing a Survey
MAVIL was implemented using SoSci Survey, a survey design and implementation software free for academic purposes [54]. The assessment was tested with four data visualizations throughout the pilots and two of them were used in the survey conduction, with each respondent always encountering two visualizations to minimize bias from any single visualization. To prevent order effects [8], automated urn-drawing method randomized the display sequence. This section discusses the rationale behind our visualization selection for testing and outlines survey tasks, which are fully detailed in the supplementary question catalog.
@@ 4.1.1 Selection of Visualization Examples on Climate Data
Data visualizations play a crucial role in scientific communication, such as when conveying critical information on the climate discourse [60, 66, 72]. While they can be effective tools for conveying complex information [65, 31], even the Intergovernmental Panel on Climate Change has acknowledged challenges in making them comprehensible to broader audiences [56, 40]. Studying their effectiveness is essential to understanding audience engagement with climate data better and ensuring trust in climate science [73].
The selection of data visualizations for MAVIL testing followed a number of criteria: they had to present climate-related data to match our topic setting, be simple enough to encourage engagement, they had to differ in chart type and remain static rather than interactive. Additionally, all visualizations were licensed under Creative Commons (CC BY-ND) for public use and credited accordingly. Their suitability for a general audience was judged during short feedback exchanges with visualization researchers.
Based on these criteria, we selected two visualizations with climate data for our survey, in particular a line chart and a bar chart (see Figure 2) were selected as they are among the most widely used visualization types [61], ensuring that respondents could engage with them regardless of prior expertise. The line chart (2(a)) shows sea level rise from 1993 to 2021, with an annotation highlighting an average annual increase of $3.5$ millimeters. The line is shaded with a light blue area labeled "Divergence," emphasizing the rate of increase. The bar chart (2(b)) compares $\text{CO}_{2}$ emissions of $10$ countries in 1990 and 2018, using color-coded bars and percentage changes. Each country’s flag is displayed and a $\text{CO}_{2}$ cloud icon is placed in the middle of the chart. During testing, both charts were in German, as this was the primary language of the surveyed audience.
@@ 4.1.2 Survey Tasks
Across all testing phases, respondents were presented with two data visualizations, meaning that Dimensions 1, 2, 3, and 4 were applied once per visualization example. Questions on Contextualizing appeared only once. Dimension 5, which addresses Reading, featured four tasks in true-or-false mode, two per visualization. Correct data reading responses across visualization examples were added to a Data Reading Score (DRS) for the analysis. An additional data reading task was designed by us and posed in reference to the line chart as an attention check to ensure survey data quality. This question had to be answered in an obligatory text field and functioned as an attention test.
To cater to a general audience, there are no prerequisites for taking the survey. Respondents are first informed about the study’s purpose of investigating the understanding of climate data visualizations. They have to confirm their consent to participate in the study and that they were not underage. We aimed to keep instructions throughout the survey concise but informative. Participants are asked to engage with two data visualizations, focusing on both content and overall impression. After viewing each visualization, they proceed to a set of related questions. Whenever a question refers directly to a visualization, the corresponding chart is displayed beneath it. Self-perceived ability ratings are presented on a five-point scale ($1$ = I disagree, $2$ = I somewhat disagree, $3$ = Undecided, $4$ = I somewhat agree, $5$ = I agree). For analysis, ratings of $1$ and $2$ were grouped to represent disagreement, and $4$ and $5$ as agreement. A rating of $3$ was considered a neutral answer. In Dimension 4 on Critiquing, respondents could provide optional text input, which was evaluated qualitatively by grouping responses into categories.
@@ 4.2 Qualitative Testing: Survey Pilots
Qualitative survey testing was carried out with eight respondents who filled out the assessment in the presence of a researcher. In semi-structured interviews, respondents described their interpretation of the two data visualization examples and discussed their experience answering MAVIL questions.
@@ 4.2.1 Purpose and Procedure of Survey Pilots
The survey pilots supported sharpening and validating MAVIL questions. By analyzing the respondents’ assessment experience, including question understanding, the feasibility of the questions could be qualitatively evaluated and then iterated upon. The insights were particularly productive for fine-tuning question phrasing. Respondents were recruited purposively through professional and personal relations, aiming for diverse educational and professional backgrounds. Six out of eight respondents had no expertise in climate data visualizations, one respondent was familiar with data visualizations due to their field of work, and one was deliberately a visualization researcher. The survey pilots were split into two phases, each consisting of four pilots. Sessions lasted between 28 and 47 minutes (median: 35 min). Interviews were audio-recorded with the consent of all participants and then transcribed.
4.2.2 Analysis of Interview Data from Survey Pilots and Learning Excerpts
The interview data was analyzed thematically by accumulating interview passages into themes and deriving a subsequent interpretation for assessment development. The analytical process took place in reference to Grounded Theory [19]. This qualitative method combines interview data into bundled theoretical interpretations justified by the subject matter. Interview data was coded and then assigned to categories [11] to gain an overview of respondents’ assessment experience. The following two themes are exemplary as categories used for further survey development, which are based on the coded interview data:
Issues of Question Comprehensibility. In the first pilot phase, all respondents pointed out unclear question phrasing, with some questions rereading. This led to specific rephrasing during the first iterative revision. In the second pilot phase, two respondents mentioned unclear phrasing, and stagnant rereading only occurred in the case of one respondent.
Engagement With Text Fields. Dimension 4 on Critiquing consisted of, among other questions, two optional text fields per visualization example. The survey pilots showed varying attitudes towards expression in text form. Three out of eight respondents principally skipped text fields; two articulated some thought; another two did not fill in the fields, even though they verbally expressed suitable inputs to the present researcher. The eighth respondent filled out all text fields conscientiously to, as was self-stated, perform well in the survey. So that no assumptions are made about the respondent’s expressive ability or their response attitude, text inputs as indicators for critical ability were kept optional. Later in this paper, the text inputs (Q4d/Q4e) are evaluated as additional insight into critical ability (in Section 5.4).
@@ 4.3 Quantitative Testing: Representative Survey
After the qualitative survey pilots, the assessment questions and the survey design were iterated to present a validated version of MAVIL to a survey panel. The panelists represented a German-speaking country’s age group distribution from 18 to 74 years and the male-female gender split within those age groups. This survey was implemented with the aid of a market research organization. Valid survey cases had to pass certain quality criteria and the resulting data set, evaluated in Section 5, held 438 cases.
4.3.1 Data Collection in Cooperation With a Market Research Organization
So that MAVIL could undergo an implementation with a representative scope, respondents were recruited in cooperation with a market research organization. The organization guided respondents to the survey in a controlled manner and pre-selected them based on crossing the agreed-upon quotas of age groups and gender. The respondents originated from the organization’s database or other websites promoting market research participation. Respondents received incentives, whose form depended on the panelist’s affiliation with either the market research organization or other platforms. The organization held demographic self-reported data from the panel, the accuracy of which was verified by collecting demographic data in our survey.
During the cooperation with the market research organization, 457 respondents completed the assessment over 14 days via an online survey. The number of completed surveys was achieved in subsets of 50 completions. We evaluated the data quality of the completed surveys after each subset, followed by the criteria for valid survey cases, as illustrated in the next subsection.
@@ 4.3.2 Quality Criteria for Valid Survey Cases
The completed survey cases were downloaded as a data set from SoSci, where the survey had been self-administrated on our part. All 457 cases from the market research cooperation period from March 3 until March 17, 2023, were selected. A completed run held responses for all obligatory questions (i.e., excluding the optional text fields), and both tested data visualizations. Respondents who exited the survey early were omitted. Another 19 cases were disregarded, as these respondents failed to answer the attention question (MAVIL Q5AT, see Figure 1) correctly. Referring to the surveyed bar chart (2(b)), this question asked respondents to name the first country listed in the $\text{CO}_{2}$ emissions chart. The correct answer was China, with variations like "china" or "Chin" also accepted.
The resulting data set, consisting of completed surveys during which the attention test was passed, shows 438 valid cases. The survey data set will be made openly accessible upon publication of this paper.
@@ 4.3.3 Demographics of the Survey
Demographic information on the respondents included their gender, age group, and education level. Respondents also had to confirm their country of residence where the survey was conducted. Inclusive options for gender self-assignment were offered. Of the 438 respondents, 223 were male ($50.9\%$) and 214 ($48.9\%$) female. One respondent reported their gender as inter. Age groups were recorded in categories of 10-year ranges and represented the distribution of the surveyed population. The minimum registered age was 18 since the legal age was a consent criterion, and the maximum was 74 years, as this was the highest age of online panelists available to the market research organization.
@@ 5 Results
MAVIL was tested in an online survey with $438$ respondents representative of a German-speaking country (age groups 18 to 74 years and male-female gender split). In the following, the results from this implementation are showcased, as they provide the first indications for the feasibility of the assessment. The survey data is analyzed to share insights into the ability composition of the surveyed population and the understanding of a line and bar chart showing climate data (see Figure 2).
@@ 5.1 Insights into Performance Measures
To give a general impression about completing the assessment: It took respondents an average of $2.75$ minutes to answer the main body of MAVIL questions per data visualization. On average, framing elements of the survey, such as the introduction and the query of demographic info, took about a minute to complete.
Based on the data reading tasks posed in MAVIL Dimension 5, there were clear deficits when evaluating the surveyed population’s Visual Data Literacy from a task-based point of view. Overall, $48\%$ of the respondents made mistakes when reading the simple line and bar chart. Effect sizes of differences between the two charts are relatively small, as seen when comparing the answer distribution of correct answers (see Figure 3). The results of the data reading tasks in Dimension 5 across the visualizations were summarized into a Data Reading Score (DRS), which corresponds to the sum of all correct data reading answers per respondent (see Table 2).
Figure 3: Data reading tasks (Q5a-Q5d) were presented in true-or-false mode (see Figure 2). In this chart, the percentages of correct and false responses given are shown, the total number of answers being 438.
Line Chart
2 Correct
Line Chart
1 Correct
Line Chart
0 Correct
Bar Chart
2 Correct
52% 18% 3%
Bar Chart
1 Correct
13% 5% 2%
Bar Chart
0 Correct
4% 3% -
Table 2: Respondents (in %) with correct answers for the data reading questions.
5.2 Respondents’ Self-Assessed Abilities and Understanding of the Surveyed Charts
In contrast to the data reading performance, a considerably smaller portion of the respondents indicate that they are not able to summarize the content of the shown data visualizations (MAVIL Q1c, see Figure 1). In the case of both charts, only $5\%$ of respondents state that they cannot summarize the visualization. This suggests the possibility of misleading information transfer, as many respondents are unaware of reading deficits. Regarding the displayed chart types (Q2c), $20\%$ of respondents did not indicate familiarity with the line chart, and $19\%$ did not indicate familiarity with the bar chart. The intersection of these statements results in $12\%$ of respondents who felt unfamiliar or uncertain with the line and the bar chart as graph types. Further, respondents have self-assessed deficits in understanding simple measurement units (Q2a). $22\%$ of respondents did not fully comprehend the usage of millimeters, and $28\%$ of respondents found millions of tons somewhat incomprehensible.

### v3

@@ 3 MdamV: A Multidimensional Assessment Method for Situated Visualization Understanding
Dimension 1: Comprehending focuses on overall perceived understanding, by addressing general impressions of the provided data visualization (Q1a/b). Further, assuming that abstract thinking is an indicator of a high level of literacy [8], respondents are asked if they could easily summarize the content of the visualization to someone (Q1c).
@@
Dim2: Decoding considers self-perceived decoding of graph specifics, such as graph elements and familiarity with the displayed chart type. This includes querying perceived comprehensibility of color usage (Q2b) and depicted measurement units (Q2a), which reflects that visualization comprehension can depend on both structural/content-related features and visual/aesthetic design elements [77]. Moreover, graph familiarity is said to be beneficial for reading data correctly [33], and therefore, is also investigated (Q2c).
@@
Dim3: Aestheticizing, we adapted the BeauVis scale [44], which consists of synonymous English-language adjectives aimed at capturing aesthetic perception. The terms were translated into German, the primary language for testing (see Sec. 4.3), yielding three items (Q3a-c). To complement these, we added a question on perceived trustworthiness (Q3d), as attributing beauty to a visualization may favorably bias its interpretation [19].
Figure 1: Overview of MdamV’s structure with six assessment dimensions (Dim1–Dim6), alongside demographic questions. For the full question catalog, see the supplementary material.
@@
Figure 2: Data visualizations used in the MdamV survey with a German-speaking audience. Figure (a) shows an official English version of the tested line chart, while Figure (b) displays the original bar chart used in the survey, which is only available in German. See Sec. 4.1.1 for details.
Dim4: Critiquing centers on critical ability and understanding of visualization construction. Respondents are asked to articulate specific feedback on the displayed data visualization. This includes indicating whether they have any erroneous findings to report (Q4a) or specific amendment proposals to share (Q4b/c), as literacy is also said to cover the ability to question an object or seek out possible changes [18] as well as to create visualizations [11, 37]. Respondents were offered subsequent optional text fields asking for visualization strengths and weaknesses (Q4d/e).
Dim5: Reading uses task-based measures in the form of data reading questions (Q5a/b), a well-established way to assess literacy performance (see Sec. 2.3). We pose true-or-false questions (see Figure 2), which directly reference the data content of the tested visualizations.
Dim6: Contextualizing covers two context-related facets of visual data literacy: perceived numeracy and topic knowledge. A higher level of numeracy has been shown to increase the likelihood of reading a data visualization correctly [76, 88], which is why we asked respondents to rate their own numeracy level (Q6b). In addition, we asked about personal interest in the visualization’s topic (Q6a), as it may influence the understanding of visually represented data [33].
Though most questions are generally phrased, some reference the data visualization example to exemplify parts of the question to respondents. This is the case for data reading tasks (Q5a/b) bound to the visualization content. Specifications were also made to Q2a by listing measurement units and Q2c by naming the chart type. In Q6a, respondents were asked to rate personal interest in the displayed topic.
@@ 4 Methodology
The design and testing of MdamV were split into four phases. At first, questions were developed in a theory-driven approach (see Sec. 3) and placed within a digital survey format. Then, in two phases of qualitative testing, the assessment was presented to eight participants. Semi-structured interviews allowed us to gain insights into the validity of MdamV questions, which were followed by iterative development. Finally, the assessment was completed by 438 respondents representing the Austrian population’s age groups, from 18 to 74 years, and their male-female gender distribution. MdamV was developed and tested in the country’s primary language. The English version presented here is a translation and has not been tested. The original questions are made available at an anonymized link (to be provided in the camera-ready version).
4.1 MdamV in Practice: Designing a Survey
MdamV was implemented using SoSci Survey, a survey design software free for academic purposes [60]. The assessment was tested with four data visualizations throughout the pilots, and two of them were used in the survey conduction, with each respondent always encountering two visualizations to minimize bias from any single visualization. To prevent order effects [9], automated urn-drawings randomized the display sequence. This section discusses the rationale behind our visualization selection for testing and outlines survey tasks, which are fully detailed in the supplementary question catalog.
@@ 4.1.1 Selection of Visualization Examples on Climate Data
As highlighted in Sec. 2.1, considering visualization-centered factors is valid when treating visual data literacy as a situated process, because each visualization can expose specific insights into how readers construct understanding. To ground our study in a relevant context, we focused on climate data visualizations, since they are an essential medium for communicating critical information on the climate discourse [66, 72, 79, 85]. While they can be effective tools for conveying complex information [71, 33], even the Intergovernmental Panel on Climate Change has acknowledged challenges in making them comprehensible to broader audiences [62, 43]. Studying their effectiveness is essential to understanding audience engagement with climate data better and ensuring trust in climate science [80].
The selection of data visualizations for MdamV testing followed a number of criteria: they had to present climate-related data to match our topic setting, be simple enough to encourage engagement, they had to differ in chart type and remain static rather than interactive. Additionally, all visualizations were licensed under Creative Commons (CC BY-ND) for public use and credited accordingly. Their suitability for a general audience was judged during short feedback exchanges with visualization researchers.
Based on these criteria, we selected a line chart and a bar chart (see Fig. 2), two of the most widely used visualization types [67, 5], to ensure that respondents could engage with them regardless of prior experience. The line chart (Fig. 2(a)) shows sea level rise from 1993 to 2021, with an annotation highlighting an average annual increase of $3.5$ millimeters. The line is shaded with a light blue area labeled "Divergence". The bar chart (Fig. 2(b)) compares $\text{CO}_{2}$ emissions of $10$ countries in 1990 and 2018, using color-coded bars and percentage changes. Each country’s flag is displayed, and a $\text{CO}_{2}$ cloud icon is placed in the middle of the chart. During testing, both charts were in German, as this was the primary language of the surveyed audience.
@@ 4.1.2 Survey Tasks
During testing, respondents were shown two data visualizations, so some assessment dimensions (Dim1-4) appeared twice, once for each visualization. Questions on Contextualizing (Dim6) appeared only once. Reading questions (Dim5) featured four tasks in true-or-false mode, two per visualization. An additional reading task for the bar chart, which had to be answered in an obligatory text field, served as an attention check to ensure survey data quality.
To cater to a general audience, there were no prerequisites for taking the survey. Respondents were first informed about the study’s purpose of investigating the understanding of climate data visualizations. They had to confirm their consent to participate in the study and that they were not underage. Participants were asked to engage with two data visualizations, focusing on both content and overall impression. After viewing each visualization, they proceeded to a set of related questions. Whenever a question referred directly to a visualization, the corresponding chart was displayed beneath it. Self-perceived ability ratings were presented on a five-point scale ($1$ = I disagree, $2$ = I somewhat disagree, $3$ = Undecided, $4$ = I somewhat agree, $5$ = I agree). For analysis, ratings of $1$ and $2$ were grouped to represent disagreement, and $4$ and $5$ as agreement. A rating of $3$ was considered a neutral answer. In Dimension 4 on Critiquing, respondents could provide optional text input, which was evaluated qualitatively by grouping responses into categories.
@@ 4.2 Qualitative Testing: Survey Pilots
The survey was piloted qualitatively with eight participants, who completed MdamV questions on two visualizations in the presence of a researcher. In semi-structured interviews, participants discussed their survey experience and explained how they interpreted the tested data visualizations.
@@ 4.2.1 Purpose and Procedure of Survey Pilots
The survey pilots helped refine and validate MdamV questions by qualitatively evaluating respondents’ experience and iterating on question clarity and feasibility. The insights were particularly productive for fine-tuning question phrasing. Respondents were recruited purposively through professional and personal relations, aiming for diverse educational and professional backgrounds. Six out of eight respondents had no expertise in climate data visualizations, one respondent was familiar with data visualizations due to their field of work, and one was a visualization researcher. The survey pilots were split into two phases, each consisting of four pilots. Sessions lasted between 28 and 47 minutes (median: 35 min). Interviews were audio-recorded with the consent of all participants and then transcribed.
4.2.2 Analysis of Survey Pilots and Learning Excerpts
The interview data was analyzed thematically by accumulating interview passages into themes and deriving a subsequent interpretation for assessment development. The analytical process took place in reference to Grounded Theory [20]. This qualitative approach combines interview data into bundled theoretical interpretations justified by the subject matter. Interview data was coded and then assigned to categories [12] to gain an overview of respondents’ experience. The following two themes are exemplary as categories used for further survey development, which are based on the coded interview data:
Issues of Question Comprehensibility. In the first pilot phase, all respondents noted unclear phrasing, with some questions requiring rereading, which led to targeted revisions. In the second phase, only two respondents mentioned unclear phrasing, and rereading only occurred in just one case.
Engagement With Text Fields. Critiquing included two optional text fields per data visualization. Participants’ willingness to use them varied, reflecting different attitudes towards expression in text form. Three out of eight respondents principally skipped text fields; two articulated some thought; another two did not fill in the fields, even though they verbally expressed suitable inputs to the present researcher. The eighth respondent filled out all text fields conscientiously, as was self-stated, to perform well in the survey. Since pilot evaluation showed that respondents varied in how they expressed themselves, to avoid bias from differences in expressive ability or response attitude, text inputs for assessing critical ability were kept optional.
@@ 4.3 Quantitative Testing: Representative Survey
After piloting, we refined MdamV’s questions and survey implementation to collect data from a panel representative of the Austrian population’s age groups (18-74 years) and male-female gender distribution within those age groups (N=$438$).
4.3.1 Data Collection With a Market Research Organization
So that MdamV could undergo an application with a representative scope, respondents were recruited in cooperation with a market research organization. The organization guided respondents to the survey in a controlled manner and pre-selected them based on crossing the agreed-upon quotas of age groups and gender. The respondents originated from the organization’s database or other websites promoting market research participation. Respondents received incentives, whose form depended on the panelist’s affiliation with either the market research organization or other platforms. The organization held demographic self-reported data from the panel, the accuracy of which was verified by collecting demographic data in our survey.
During a period of two weeks, respondents were directed to our online survey, with survey cases achieved in subsets of 50 completions. We evaluated the data quality of survey completions after each subset, followed by the criteria for valid survey cases, as illustrated in the next subsection.
@@ 4.3.2 Quality Criteria for Valid Survey Cases
The completed survey cases were downloaded as a data set from SoSci, where the survey had been self-administered on our part. All 457 cases from the market research cooperation period from March 3 until March 17, 2023, were selected. A completed case held responses for all obligatory questions (i.e., excluding the optional text fields), and both tested data visualizations. Respondents who exited the survey early were omitted. Another 19 cases were disregarded, as these respondents failed to answer the attention question (see Q5AT in Fig. 1) correctly. Referring to the surveyed bar chart (Fig. 2(b)), this question asked respondents to name the first country listed in the $\text{CO}_{2}$ emissions chart. The correct answer was China, with variations like "china" or "Chin" also accepted. The resulting data set, with 438 cases that passed the attention test, will be made openly accessible upon publication of this paper.
@@ 4.3.3 Demographics of the Survey
Demographic information on the respondents included their gender, age group, and education level. Respondents had to confirm their country of residence where the survey was conducted. Inclusive options for gender self-assignment were offered. Of the $438$ respondents, $223$ were male ($50.9\%$) and 214 ($48.9\%$) female. One respondent reported their gender as inter. Age groups were recorded in categories of 10-year ranges and represented the distribution of the surveyed population. The minimum registered age was 18 since the legal age was a consent criterion, and the maximum was 74 years, as this was the highest age of online panelists available to the market research organization.
@@ 5 Results
We tested MdamV with $438$ respondents, representative of Austria’s population in terms of age groups (18-74 years) and male-female gender split. Using MdamV on a line and a bar chart of climate data (see Fig. 2), we gathered insights into respondents’ visualization understanding. The following results illustrate the feasibility of our proposed method for assessing situated visualization understanding.
Figure 3: Data reading tasks (Q5a-Q5d) were presented in true-or-false mode (see Fig. 2). (a) shows correct vs. false responses and (b) the distribution of correct answers per chart, both with N=438.
@@ 5.1 Insights into Performance Measures
To give a general impression about completing the assessment: It took respondents an average of $2.75$ minutes to answer the main body of MdamV questions per data visualization. On average, framing elements of the survey, such as the introduction and the query of demographic info, took about a minute to complete.
Based on the Reading questions, there were clear deficits when evaluating the surveyed population’s visualization understanding from a task-based point of view. Overall, $48\%$ of the respondents made mistakes when reading the simple line and bar chart. Effect sizes of differences between the two charts are relatively small, as seen when comparing the answer distribution of correct answers (see Fig. 3a). Just over half of respondents solved all four data reading tasks correctly, while none answered all of them incorrectly (see Fig. 3b).
5.2 Respondents’ Perception of the Surveyed Charts
In contrast to the data reading performance, a considerably smaller portion of the respondents indicated that they are not able to summarize the content of the shown data visualizations (see Q1c in Fig. 1). In the case of both charts, only $5\%$ of respondents stated that they cannot summarize the visualization. This suggests the possibility of misleading information transfer, as many respondents are unaware of reading deficits. Regarding the displayed chart types (Q2c), The intersection of these statements results in $12\%$ of respondents who felt unfamiliar or uncertain with the line and the bar chart as chart types. Further, respondents have self-assessed deficits in understanding simple measurement units (Q2a). $22\%$ of respondents did not fully comprehend the usage of millimeters, and $28\%$ of respondents found millions of tons somewhat incomprehensible.